Circular No. 6839

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1997 XF_{11}

E. F. Helin, Jet Propulsion Laboratory, reports that K. J. Lawrence has located prediscovery observations of this object in 1990 on films taken by Helin, Lawrence and B. Roman in the course of the Planet-Crossing Asteroid Survey with the 0.46-m Schmidt telescope at Palomar:

1990 UT  α_{2000}  δ_{2000}  B
Mar. 22.44427  13^{h}15^{m}16.^{s}50  −12^{o}54′21′′.2  17.3
22.47674  13 15 13.91  −12 54 04.3
23.44062  13 14 00.86  −12 45 47.4
23.46632  13 13 58.73  −12 45 34.4

The improved orbital elements below have been computed by G. V. Williams, Harvard-Smithsonian Center for Astrophysics, from 99 observations at the 1990 and 1997-1998 apparitions, the mean residual being 0′′.58:

Epoch = 1997 Dec. 18.0 TT

T = 1997 July 1.1954 TT  ω = 102.^{o}4645
e = 0.483775  Ω = 214.1319 }2000.0
q = 0.744247 AU  i = 4.0948

a = 1.441710 AU n^{o} = 0.5693602 P = 1.731 years

1998 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Mar. 8 6^{h}15.^{m}92 +17^{o}56.^{′}7 1.557 2.070 106.^{o}5 27.^{o}4 20.8
13 6 19.04 +18 08.0 1.629 2.080 102.3 27.8 20.9
18 6 22.81 +18 18.0 1.701 2.089  98.2 28.1 21.0
23 6 27.18 +18 26.5 1.773 2.098  94.3 28.3 21.1
28 6 32.08 +18 33.4 1.846 2.105  90.5 28.3 21.2
Apr. 2 6 37.46 +18 38.7 1.917 2.112  86.8 28.2 21.3
7 6 43.26 +18 42.1 1.988 2.118  83.3 28.0 21.4

This new computation shows that the encounter with the earth in 2028 will occur on Oct. 26.3 UT, the miss distance then being 0.0064 AU. A computation from the 60-day-arc orbit on MPC 31283 indicated a comparable distance (0.0055 AU). With regard to the remarks on IAUC 6837, it should be noted that, from the 88-day arc of data then available, several independent authorities gave comparable nominal miss distances for 2028 Oct. 26: D. K. Yeomans and P. Chodas, Jet Propulsion Laboratory, derived 0.00058 AU; K. Muinonen, University of Helsinki, 0.00033 AU; and E. Bowell, Lowell Observatory, 0.00023 AU. The new computation shows that there were also approaches to 0.032 AU in 1971 and to 0.015 AU in 1957.

1998 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden